OpenStudy (anonymous):

f(x)=5x((5-5x)^(1/2)) find the equation of line tangent to the graph f at x=-2

myininaya (myininaya):

so have you found f'(x)?

OpenStudy (anonymous):

don't think it's the right one

myininaya (myininaya):

\[f'(x)=5[1*(5-5x)^\frac{1}{2}+x*\frac{1}{2}(5-5x)^{\frac{1}{2}-1}(-5)]\]

myininaya (myininaya):

\[f'(x)=5*\sqrt{5-5x}-\frac{5}{2\sqrt{5-5x}}=\frac{5\sqrt{5-5x}*2\sqrt{5-5x}-5}{2\sqrt{5-5x}}\] =\[\frac{10(5-5x)-5}{2\sqrt{5-5x}}=\frac{50-50x-5}{2\sqrt{5-5x}}\] =\[\frac{-5x+45}{2\sqrt{5-5x}}*\frac{\sqrt{5-5x}}{\sqrt{5-5x}}=\frac{(-5x+45)\sqrt{5-5x}}{2(5-5x)}\] =\[\frac{-5(x-9)\sqrt{5-5x}}{2(5)(1-1x)}=\frac{-(x-9)\sqrt{5-5x}}{2(1-x)}\]

myininaya (myininaya):

i was just trying to make it look nicer

myininaya (myininaya):

anyways to find the slope find f'(-2)

OpenStudy (anonymous):

so I plug it in an my x value at -2 is 7.1005. I'm stuck after that.

myininaya (myininaya):

\[f'(-2)=\frac{-(-2-9)\sqrt{5-5(-2)}}{2(1-(-2))}=\frac{11\sqrt{5+10}}{6}=\frac{11\sqrt{15}}{6}\]

myininaya (myininaya):

do we know a point in the tangent line at x=-2? yes we (-2,f(-2)) on on that line \[f(-2)=5(-2)\sqrt{5-5(-2)}=-10\sqrt{15}\] a line has the form y=mx+b m=11sqrt{15}/6 (x1,y1)=(-2,-10sqrt{15}) \[-10\sqrt{15}=\frac{11\sqrt{15}}{6}*(-2)+b\] find b to find the y-intercept of the tangent line

myininaya (myininaya):

\[b=-10\sqrt{15}+\frac{11\sqrt{15}*2}{6}\] \[b=\sqrt{15}(-10+\frac{11*2}{6})=\sqrt{15}(-10+\frac{11}{3})=\sqrt{15}(\frac{-30+11}{3})\] \[b=\frac{-19}{3}\sqrt{15}\]

OpenStudy (anonymous):

I get 7.100469468x+52.9307724, which is wrong. What am I doing?

myininaya (myininaya):

\[y=\frac{11\sqrt{15}}{6}*x+\frac{-19\sqrt{15}}{3}\]

myininaya (myininaya):

first does say to find the exact line or an appoximation also are y-intercepts are way different

OpenStudy (anonymous):

all it says is find the equation of line tangent to the graph, and tangent line : y=?

OpenStudy (anonymous):

I plugged it in, still not getting it

myininaya (myininaya):

line has form y=mx+b we know a point on the line and we know the slope so we can find the y-intercept

myininaya (myininaya):

what did you plug in for the point?

OpenStudy (anonymous):

exactly what you gave me, and others that haven't worked either.

myininaya (myininaya):

how did you get a different y-intercept then?

OpenStudy (anonymous):

incorrect math, I guess.

OpenStudy (anonymous):

Good work myininaya. These computer exercise would reject an answer for little mistakes. There is a little mistake at the derivative. Find derivative of inner function (5-5x); the derivative is -5

myininaya (myininaya):

yeah i have that above

OpenStudy (anonymous):

I give up.

myininaya (myininaya):

show me what you did when you tried to find the y-intercept

myininaya (myininaya):

you said you plug in exactly what i gave right and got different y-intecept than me right? y=mx+b b=y-mx b=f(-2)-f'(2)(-2)

myininaya (myininaya):

b=f(-2)+2f'(2)

myininaya (myininaya):

b=-10sqrt{15}+2(11sqrt{15}/6)

myininaya (myininaya):

you should get a negative y-intercept since 10>11/3

myininaya (myininaya):

b=sqrt{15}*(-10+11/3) b=sqrt{15}*(-30/3+11/3) b=sqrt{15}*(-19/3)

OpenStudy (anonymous):

Try this y=25.8207x + 12.9099

OpenStudy (anonymous):

chaguanas, that worked :] however did you find it?

myininaya (myininaya):

i think i forgot to distrubute the 5 earlier when i was doing f'(x)

OpenStudy (anonymous):

I used the KISS method: Keep it simple. :)

OpenStudy (anonymous):

I had no idea what I was doing wrong. Do you think you could help me with thesame type of problem? f(x)= (3x^3)/((4-5x)^5), equation of line tangent to graph of f at x=2.

OpenStudy (anonymous):

Myininaya, you were doing it right, but this is one of those computerized thingies. You made a lot of transpositions and transformations to 'make it look pretty' and you might have threw it off one percentage point. Let's see this one.

myininaya (myininaya):

i did notice some errors above

myininaya (myininaya):

i forgot to distribute something and then forgot to add something lol

OpenStudy (anonymous):

\[f'(x)=[9x ^{2}(4-5x)^{5}+75x ^{3}(4-5x)^{4}]/(4-5x)^{10}\]

OpenStudy (anonymous):

I get \[f \prime(x)= 6x ^{2}(5x+6)/(4-5x)^6\]

OpenStudy (anonymous):

with that, I get m=.00823

myininaya (myininaya):

y=mx+b -1/324=(2/243)*2+b b=-1/324-4/243 b=-19/972 y=2/243*x-19/972 but maybe i made a mistake again lol

OpenStudy (anonymous):

Show us how you got that f'(x)

myininaya (myininaya):

your f' looks delicious gj

myininaya (myininaya):

did you look at the attachment yet?

OpenStudy (anonymous):

That was for mnad1, looks like he made mistakes in derivative. I have an old computer and I can't read a lot of the attachments.

OpenStudy (anonymous):

your answer is correct, myininaya. Thank you! I'm annoyed because I get a completely different derivative. Looking at your attachment I see that I deviate from the correct steps pretty much at the beginning

OpenStudy (anonymous):

but you guys are awesome, thanks for the elaborate explanations, it really helped

myininaya (myininaya):

you got 6x^2(5x+6))/(4-5x)^6 right?

myininaya (myininaya):

ok well your welcome lol

OpenStudy (anonymous):

OK, I keep telling him that is wrong. How did you get that myininaya, compare it with my derivative about five messages up.

myininaya (myininaya):

yours is right too, i just factored (4-5x)^4 out on top and canceled (4-5x)^4 on top and bottom leaving me with on top: 9x^2(4-5x)+75x^3 on bottom: (4-5x)^6 we have like terms on top:36x^2-45x^3+75x^3=36x^2+30x^3=6x^2(6+5x)

OpenStudy (anonymous):

yeah, that was my f'(x). His was the same as yours, and you were evidently both right.

myininaya (myininaya):

hers*

myininaya (myininaya):

i was just going for pretty looks again lol

OpenStudy (anonymous):

OK, mnad1, sorry I keep saying you're wrong. But in a case like this you are inputting values of x, so there is no need to break it down further. It increases the chance of making a mistake. Only a pro like myininaya could do that.

myininaya (myininaya):

lol

myininaya (myininaya):

but as we learned i can make mistakes too example the first one we did

myininaya (myininaya):

yes simplifying things can be lead to problems like mistakes so if your teacher doesn't care about simplying (or making pretty) then just do what chag says well unless you just enjoy it like i do

myininaya (myininaya):

but mnad the form we have it in we can easily find horiztonal tangents :)

OpenStudy (anonymous):

I guess you could say I'm half assing it cause in order to get the x value, I just plug the derivative in my calc and look at the table. It's not a problem unless the derivative is wrong, as it was in my case and I had no idea what my error was. until now.

myininaya (myininaya):

we can say we have a horizontal tangent at x=0 and x=-6/5 if we were asked to find this

OpenStudy (anonymous):

I don't know if you noticed, you might have, but math is not my best class alas.

myininaya (myininaya):

i didn't notice

myininaya (myininaya):

be positive and don't give up if you believe you can't do, then chances are you probably won't be able to do it. so just believe you can

OpenStudy (anonymous):

it's a summer class so we're going through the material at a faster pace, so I have to apply extra effort to learn it. I'm glad I came across OS though, the people here are extremely helpful.